OpenStudy (anonymous):

Find g(1) if g(x) = x 2 + 1.

hero (hero):

\[g(x) = x^2 + 1\] \[g(1) = 1^2 + 1\]

hero (hero):

Simplify \(g(1)\)

OpenStudy (anonymous):

ooooooo ok

hero (hero):

Notice that to solve \(g(1)\), you set \(x = 1\)

hero (hero):

Then simplify

OpenStudy (anonymous):

thanks for your help i got it from here